# Model Reconstruction Using Counterfactual Explanations: Mitigating the Decision Boundary Shift

### v1

@@ 3 Main Results
3.1 Fundamental guarantees on model extraction using closest counterfactuals
We provide a theoretical guarantee on model extraction using the closest counterfactuals in Theorem 3.2. We first start out with demonstrating how the closest counterfactuals provide a linear approximation of any decision boundary. Prior work (Yadav, Moshkovitz, and Chaudhuri 2023) shows that for linear models, the line joining a query instance $\boldsymbol{x}$ and the closest counterfactual $\boldsymbol{w}(=g_{m}(\boldsymbol{x}))$ is perpendicular to the linear decision boundary. We generalize this observation to any differentiable decision boundary, not necessarily linear, as follows:
@@ 3.1 Fundamental guarantees on model extraction using closest counterfactuals
Figure 3: Polytope approximation of a convex decision boundary.
The proof follows by showing that the $d$-dimensional ball with radius $||\boldsymbol{x}-\boldsymbol{w}||_{2}$ touches (as in Definition 2.5) $\mathcal{S}$ at $\boldsymbol{w}$, and invoking Lemma 2.6. For details see Appendix A.1.
As a direct consequence of Lemma 3.1, an adversary may query the system and calculate tangent hyperplanes of the decision boundary drawn at the counterfactual instances. This leads to a linear approximation of the decision boundary (see Fig. 3).
If the decision boundary is convex from the unfavorable side, i.e., the region where $\left\lfloor{m(\boldsymbol{x})}\right\rceil=0$, such an approximation will provide a set of supporting hyperplanes. The intersection of these supporting hyperplanes will provide a circumscribing polytope approximation of the decision boundary. We show that the fidelity of such an approximation, evaluated over uniformly distributed input instances, tends to 1 when the number of counterfactuals is large.
@@
Let $m$ be the target binary classifier whose decision boundary is convex (i.e., the set $\{\boldsymbol{x}\in[0,1]^{d}:\left\lfloor{m(\boldsymbol{x})}\right\rceil=1\}$ is convex) and has a continuous second derivative. Denote by $\tilde{M}_{n}$, the polytope approximation of $m$ constructed with $n$ supporting hyperplanes obtained through i.i.d. queries. Assume that the fidelity is evaluated with respect to ${\mathbb{D}}_{\text{ref}}$ which is uniformly distributed over $[0,1]^{d}$. Then, when $n\rightarrow\infty$ the expected fidelity of $\tilde{M}_{n}$ with respect to $m$ is given by
@@
The proof utilizes a result from polytope theory (Böröczky Jr and Reitzner 2004) which provides a complexity bound on volume-approximating smooth convex sets by random polytopes. The proof involves observing that the volume of the overlapping decision regions of $m$ and $\tilde{M}_{n}$ (for example, regions A and C in Fig. 3) translates to the expected fidelity when evaluated under a uniformly distributed ${\mathbb{D}}_{\text{ref}}$. Appendix A.2 provides the detailed steps.
Figure 4: Approximating concave regions need denser queries than convex regions.
@@
Even though Theorem 3.2 assumes convexity of the decision boundary for analytical tractability, the attack can also be extended to a concave decision boundary. This is because the closest counterfactual will always lead to a tangent hyperplane irrespective of convexity and now the rejected region can be seen as the intersection of these half-spaces (Lemma 3.1 does not assume convexity). However, it is worth noting that approximating a concave decision boundary is, in general, more difficult than approximating a convex region. To obtain equally-spaced out tangent hyperplanes on the decision boundary, a concave region will require a much denser set of query points (see Fig. 4) due to the inverse effect of length contraction discussed in Aleksandrov (1967, Chapter III Lemma 2). Furthermore, approximating a decision boundary which is neither convex nor concave is much more challenging as the decision regions can no longer be approximated as intersections of half-spaces.
We note that the aforementioned attack relies on the closest counterfactual which is often challenging to generate. Practical counterfactual generation mechanisms often provide counterfactuals that are reasonably close but may not be exactly the closest. This motivates us to propose a more general strategy – that we call Counterfactual Clamping Attack (CCA) – that only leverages the fact that the counterfactuals lie closer to the decision boundary but need not be the closest. Experiments indicate that the query complexity of CCA is upper-bounded by the fundamental complexity bound in Theorem 3.2 (see Fig. 14 in Appendix B.2).
3.2 An attack based on the proximity of counterfactuals to the target decision boundary
Before introducing our proposed attack CCA, we first discuss our rationale behind devising this attack. Let us consider the difference of the model output probabilities (before thresholding) as a measure of similarity between the target and surrogate models because it would force the decision boundaries to overlap. Our proposed attack CCA is motivated from the following observation: the difference of two models’ output probabilities corresponding to a given input instance $\boldsymbol{x}$ can be bounded by having another point with matching outputs in the affinity of the instance $\boldsymbol{x}$.
@@ 3.2 An attack based on the proximity of counterfactuals to the target decision boundary
Suppose the target $m(\boldsymbol{x})$ and the surrogate $\tilde{m}(\boldsymbol{x})$ are Lipschitz continuous with constants $\gamma_{1}$ and $\gamma_{2}$, respectively. Assume $m(\boldsymbol{w})=\tilde{m}(\boldsymbol{w})$ for some $\boldsymbol{w}\in[0,1]^{d}$. Then, for any $\boldsymbol{x}\in[0,1]^{d}$, the difference between the outputs of the two models is bounded from above as follows: $|\tilde{m}(\boldsymbol{x})-m(\boldsymbol{x})|\leq(\gamma_{1}+\gamma_{2})||\boldsymbol{x}-\boldsymbol{w}||_{2}.$
The proof is presented in Appendix A.3. Usually, a smaller Lipschitz constant is indicative of a higher generalizability of a model (Gouk et al. 2021; Pauli et al. 2021).
Remark 3.5 (Local Lipschitz Continuity).
Global Lipschitz continuity is not always necessary. If a well-spread set of points ${\mathbb{W}}=\{\boldsymbol{w}_{i},i=1,\dots,N\}$ satisfying $m(\boldsymbol{w}_{i})=\tilde{m}(\boldsymbol{w}_{i})$ for $i=1,\dots,N$ is available over some region ${\mathbb{A}}\subseteq[0,1]^{d}$, a more relaxed condition of local Lipschitz continuity of $m$ and $\tilde{m}$ in the locality of $\boldsymbol{w}_{i}$’s is sufficient to ensure the conditions of Theorem 3.4 for all $\boldsymbol{x}\in{\mathbb{A}}$
@@
Proposed Counterfactual Clamping Attack: Theorem 3.4 provides the motivation for a novel model extraction strategy. Let $\boldsymbol{w}$ be a counterfactual. Recall that $\partial{\mathbb{M}}$ denotes the decision boundary of $m$. As implied by the theorem, for any $\boldsymbol{x}\in\partial{\mathbb{M}}$, the deviation of the surrogate model output from the target model output can be bounded above by $(\gamma_{1}+\gamma_{2})||\boldsymbol{x}-\boldsymbol{w}||_{2}$ given that all the counterfactuals satisfy $m(\boldsymbol{w})=\tilde{m}(\boldsymbol{w})$. Knowing that $m(\boldsymbol{w})=0.5$, we may design a loss function which clamps $\tilde{m}(\boldsymbol{w})$ to be 0.5. Consequently, with a sufficient number of well-spaced counterfactuals to cover $\partial{\mathbb{M}}$, we may achieve arbitrarily small $|\tilde{m}(\boldsymbol{x})-m(\boldsymbol{x})|$ at the decision boundary of $m$ (Fig. 5).
Figure 5: Rationale for Counterfactual Clamping Attack.
@@
However, we note that simply forcing $\tilde{m}(\boldsymbol{w})$ to be exactly equal to 0.5 is quite unstable since in practice the target model’s output $m(\boldsymbol{w})$ is known to be close to 0.5, while being greater but may not be exactly equal. Thus, we propose a unique loss function for training the surrogate neural networks that does not penalize the counterfactuals that are already inside the decision boundary of the surrogate model.
For $0<k\leq 1$, we define the loss function as
$L_{k}$ $(\tilde{m},y)=\frac{1}{\left|{\mathbb{D}}\right|}\sum_{\boldsymbol{x}\in{\mathbb{D}}}\Bigg{(}\mathds{1}\left[y(\boldsymbol{x})=0.5,\tilde{m}(\boldsymbol{x})\leq k\right]$
×$\times\Bigg{\{}k\log\left(\frac{k}{\tilde{m}(\boldsymbol{x})}\right)+(1-k)\log\left(\frac{1-k}{1-\tilde{m}(\boldsymbol{x})}\right)\Bigg{\}}$
$\hskip 85.35826pt+\mathds{1}\left[y(\boldsymbol{x})\neq 0.5\right]L(\tilde{m},y)\Bigg{)}.$ (4)
@@
Here, $y(\boldsymbol{x})$ denotes the label assigned to the input instance $\boldsymbol{x}$ by the target model, received from the API and $L(\tilde{m},y)$ is the binary cross-entropy loss. We assume that the counterfactuals are distinguishable from the ordinary instances, and assign them a label of $y(\boldsymbol{w})=0.5$. Here, ${\mathbb{D}}$ denotes the set of all query instances and the counterfactuals. The first term accounts for the counterfactuals, where they are assigned a non-zero loss if the surrogate model’s prediction is below $k$. This term ensures $\tilde{m}(\boldsymbol{w})\approx k$ for the counterfactuals $\boldsymbol{w}$ but allows for counterfactuals that lie farther inside the favorable region, which are the result of imperfections in the generating mechanisms. The second term is the ordinary binary cross-entropy loss, which becomes non-zero only for ordinary query instances. Note that substituting $k=1$ in $L_{k}(\tilde{m},y)$ yields the ordinary binary cross entropy loss. Algorithm 1 summarizes the attack proposed above.
Algorithm 1 Proposed Counterfactual Clamping Attack
@@
3:     Query API with $\boldsymbol{x}$ to get $y(\boldsymbol{x})$$y(\boldsymbol{x})\in\{0,1\}$}
4:     ${\mathbb{A}}\leftarrow{\mathbb{A}}\cup\{(\boldsymbol{x},y(\boldsymbol{x}))\}$
5:     if $y(\boldsymbol{x})=0$ then
@@
10:  Train $\tilde{m}$ on ${\mathbb{A}}$ with $L_{k}(\tilde{m},y)$ as the loss
@@
Remark 3.6 (Extensions to Other Contours).
This strategy can be extended to approximate any given contour $\partial{\mathbb{M}}_{k}=\{\boldsymbol{x}\in[0,1]^{d}:m(\boldsymbol{x})=k\}$. It can be helpful in cases where the counterfactual generating algorithm provides counterfactuals that lie on $\partial{\mathbb{M}}_{k}$ for some $k>0.5$, e.g., due to constraints imposed by robustness (Upadhyay, Joshi, and Lakkaraju 2021; Hamman et al. 2023; Black, Wang, and Fredrikson 2022), given that the attacker knows the value of $k$. When available, the probability predictions received through the API (Tramèr et al. 2016) can be used to fine-tune $k$.
It is noteworthy that this approach is different from the broad area of soft-label learning (Nguyen, Valizadegan, and Hauskrecht 2011; Nguyen et al. 2011) in two major aspects: (i) the labels in our problem do not smoothly span the interval [0,1] – instead they are either 0, 1 or 0.5; (ii) labels of counterfactuals do not indicate a class probability; even though a label of 0.5 is assigned, there can be counterfactuals that are well within the surrogate decision boundary which should not be penalized. Nonetheless, we also perform ablation studies where we compare the performance of our proposed loss function with another potential loss which simply forces $\tilde{m}(\boldsymbol{w})$ to be exactly $0.5$, inspired from soft-label learning (see Appendix B.2 for results).
CCA overcomes two challenges beset in existing works; (i) the problem of decision boundary shift (particularly with one-sided counterfactuals) present in the method suggested by Aïvodji, Bolot, and Gambs (2020) and (ii) the need for counterfactuals from both sides of the decision boundary in the methods of Aïvodji, Bolot, and Gambs (2020) and Wang, Qian, and Miao (2022).
Figure 6: A 2-dimensional demonstration of the proposed attack. Orange and blue shades denote the favorable and unfavorable decision regions of each model. Circles denote the target model’s training data.
Table 1: Fidelity achieved with 400 queries on the real-world datasets. Values presented are the averages $\pm$ standard deviations (as percentages) over an ensemble of size 100. Target model has hidden layers with neurons (20,10). Model 0 is similar to the target model in architecture. Model 1 has hidden layers with neurons (20, 10, 5).

### v2

@@ 3 Main Results
We provide a set of theoretical guarantees on query size of model reconstruction that progressively build on top of each other, starting from stronger guarantees in a somewhat restricted setting towards more approximate guarantees in broadly applicable settings. We discuss important intuitions that can be inferred from the results. Finally, we propose a model reconstruction strategy that is executable in practice, which is empirically evaluated in Section 4.
3.1 Convex decision boundaries and closest counterfactuals
We start out with demonstrating how the closest counterfactuals provide a linear approximation of any decision boundary, around the corresponding counterfactual. Prior work (Yadav et al., 2023) shows that for linear models, the line joining a query instance $\boldsymbol{x}$ and the closest counterfactual $\boldsymbol{w}(=g_{m}(\boldsymbol{x}))$ is perpendicular to the linear decision boundary. We generalize this observation to any differentiable decision boundary, not necessarily linear.
@@ 3.1 Convex decision boundaries and closest counterfactuals
Figure 3: Polytope approximation of a convex decision boundary using the closest counterfactuals.
The proof follows by showing that the $d$-dimensional ball with radius $||\boldsymbol{x}-\boldsymbol{w}||_{2}$ touches (as in Definition2.6) $\mathcal{S}$ at $\boldsymbol{w}$, and invoking Lemma 2.7. For details see Appendix A.1. As a direct consequence of Lemma 3.1, a user may query the system and calculate tangent hyperplanes of the decision boundary drawn at the closest counterfactuals. This leads to a linear approximation of the decision boundary at the closest counterfactuals (see Fig. 3).
If the decision boundary is convex, such an approximation will provide a set of supporting hyperplanes. The intersection of these supporting hyperplanes will provide a circumscribing convex polytope approximation of the decision boundary. We show that the average fidelity of such an approximation, evaluated over uniformly distributed input instances, tends to 1 when the number of queries is large.
@@
Let $m$ be the target binary classifier whose decision boundary is convex (i.e., the set $\{\boldsymbol{x}\in[0,1]^{d}:\left\lfloor{m(\boldsymbol{x})}\right\rceil=1\}$ is convex) and has a continuous second derivative. Denote by $\tilde{M}_{n}$, the convex polytope approximation of $m$ constructed with $n$ supporting hyperplanes obtained through i.i.d. counterfactual queries. Assume that the fidelity is evaluated with respect to ${\mathbb{D}}_{\text{ref}}$ which is uniformly distributed over $[0,1]^{d}$. Then, when $n\rightarrow\infty$ the expected fidelity of $\tilde{M}_{n}$ with respect to $m$ is given by
@@
Theorem 3.2 provides an exact relationship between the expected fidelity and number of queries. The proof utilizes a result from random polytope theory (Böröczky Jr and Reitzner, 2004) which provides a complexity bound on volume-approximating smooth convex sets by convex polytopes. The proof involves observing that the volume of the overlapping decision regions of $m$ and $\tilde{M}_{n}$ (for example, regions A and C in Fig. 3) translates to the expected fidelity when evaluated under a uniformly distributed ${\mathbb{D}}_{\text{ref}}$. Appendix A.2 provides the detailed steps.
Figure 4: Approximating a concave region needs denser queries w.r.t. a convex region.
@@
This strategy of linear approximations can also be extended to a concave decision boundary since the closest counterfactual will always lead to a tangent hyperplane irrespective of convexity. Now the rejected region can be seen as the intersection of these half-spaces (Lemma 3.1 does not assume convexity). However, it is worth noting that approximating a concave decision boundary is, in general, more difficult than approximating a convex region. To obtain equally-spaced out tangent hyperplanes on the decision boundary, a concave region will require a much denser set of query points (see Fig. 4) due to the inverse effect of length contraction discussed in Aleksandrov (1967, Chapter III Lemma 2). Deriving similar theoretical guarantees for a decision boundary which is neither convex nor concave is much more challenging as the decision regions can no longer be approximated as intersections of half-spaces. The assumption of convex decision boundaries may only be satisfied under limited scenarios such as input-convex neural networks (Amos et al., 2017). However, we observe that this limitation can be circumvented in case of ReLU networks to arrive at a probabilistic guarantee as discussed next.
3.2 ReLU networks and closest counterfactuals
Rectified Linear Units (ReLU)are one of the most used activation functions in neural networks (Zeiler et al., 2013; Maas et al., 2013; Ronneberger et al., 2015; He et al., 2016). A deep neural network that uses ReLU activations can be represented as a Continuous Piece-Wise Linear (CPWL)function (Chen et al., 2022; Hanin and Rolnick, 2019). A CPWL function comprises of a union of linear functions over a partition of the domain. Definition 3.4 below provides a precise characterization.
Definition 3.4 (Continuous Piece-Wise Linear (CPWL)Function (Chen et al., 2022)).
A function $\ell:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is said to be continuous piece-wise linear if and only if
  1. 1.
    There exists a finite set of closed subsets of $\mathbb{R}^{d}$, denoted as $\{{\mathbb{U}}_{i}\}_{i=1,2,\dots,q}$ such that $\cup_{i=1}^{q}{\mathbb{U}}_{i}={\mathbb{R}}^{n}$
  2. 2.
    $\ell(\boldsymbol{x})$ is affine over each ${\mathbb{U}}_{i}$ i.e., over each ${\mathbb{U}}_{i},\ell(\boldsymbol{x})=\ell_{i}(x)=\boldsymbol{a}_{i}^{T}\boldsymbol{x}+b_{i}$ with $\boldsymbol{a}_{i}\in{\mathbb{R}}^{d}$, $b_{i}\in{\mathbb{R}}$.
This definition can be readily applied to the models of our interest, of which the domain is the unit hypercube $[0,1]^{d}$. A neural network with ReLU activations can be used as a classifier by appending a Sigmoid activation $\sigma(z)=\frac{1}{1+e^{-z}}$ to the final output. We denote such a classifier by $m(\boldsymbol{x})=\sigma(\ell(\boldsymbol{x}))$ where $\ell(\boldsymbol{x})$ is CPWL. It has been observed that the number of linear pieces $q$ of a trained ReLU network is generally way below the theoretically allowed maximum (Hanin and Rolnick, 2019).
We first show that the decision boundaries of such CPWL functions are collections of polytopes (not necessarily convex). The proof of Lemma 3.5 is deferred to Appendix A.3.
Lemma 3.5.
Let $m(\boldsymbol{x})=\sigma(\ell(\boldsymbol{x}))$ be a ReLU classifier, where $\ell(\boldsymbol{x})$ is CPWL and $\sigma(.)$ is the Sigmoid function. Then, the decision boundary $\partial{\mathbb{M}}=\{\boldsymbol{x}\in[0,1]^{d}:m(\boldsymbol{x})=0.5\}$ is a collection of (possibly non-convex) polytopes in $[0,1]^{d}$, when considered along with the boundaries of the unit hypercube.
Next we will analyze the probability of successful model reconstruction using counterfactuals. Consider a uniform grid $\mathcal{N}_{\epsilon}$ over the unit hypercube $[0,1]^{d}$, where each cell is a smaller hypercube with side length $\epsilon$ (see Fig. 5). For this analysis, we make an assumption: If a cell contains a part of the decision boundary, then that part is completely linear (affine) within that small cell  [^2] This is violated only for the cells containing parts of the edges of the decision boundary. However, we may assume that $\epsilon$ is small enough so that the total number of such cells is negligible compared to the total cells containing the decision boundary. Generally, $\epsilon$ can be made sufficiently small such that for most of the cells, if a cell contains a part of the decision boundary then that part is affine within the cell. .
Figure 5: $\mathcal{N}_{\epsilon}$ grid and inverse counterfactual regions. Thick solid lines indicate the decision boundary pieces (${\mathbb{H}}_{i}$’s). White color depicts the accepted region. Pale-colored are the inverse counterfactual regions of the ${\mathbb{H}}_{i}$’s with the matching color. In this case $k(\epsilon)=7$ and $v^{*}(\epsilon)$ is the area of lower amber region.
Now, as the decision boundary becomes linear for each small cell that it passes through, having just one closest counterfactual in each such cell is sufficient to get the decision boundary in that cell (recall Lemma 3.1). We formalize this intuition in Theorem 3.6 to obtain a probabilistic guarantee on the success of model reconstruction. A proof is presented in Appendix A.4.
@@ 3.2 ReLU networks and closest counterfactuals
Let $m$ be a target binary classifier with ReLU activations. Let $k(\epsilon)$ be the number of cells through which the decision boundary passes. Define $\{{\mathbb{H}}_{i}\}_{i=1,\dots,k(\epsilon)}$ to be the collection of affine pieces of the decision boundary within each decision boundary cell where each ${\mathbb{H}}_{i}$ is an open set. Let $v_{i}(\epsilon)=V({\mathbb{G}}_{m,g_{m}}({\mathbb{H}}_{i}))$ where $V(.)$ is the $d-$dimensional volume (i.e., the Lebesgue measure) and ${\mathbb{G}}_{m,g_{m}}(.)$ is the inverse counterfactual region w.r.t. $m$ and the closest counterfactual generator $g_{m}$. Then the probability of successful reconstruction with counterfactual queries distributed uniformly over $[0,1]^{d}$ is lower-bounded as
${\mathbb{P}}\left[\text{Reconstruction}\right]\geq 1-k(\epsilon)(1-v^{*}(\epsilon))^{n}$ (2)
where $v^{*}(\epsilon)=\min_{i=1,\dots,k(\epsilon)}v_{i}(\epsilon)$ and $n$ is the number of queries.
Remark 3.7.
Here $k(\epsilon)$ and $v^{*}(\epsilon)$ depend only on the nature of the model being reconstructed and are independent of the number of queries $n$. The value of $k(\epsilon)$ roughly grows with the surface area of the decision boundary (e.g., length when input is 2D), showing that models with more convoluted decision boundaries might need more queries for reconstruction. Generally, $k(\epsilon)$ lies within the interval $\frac{A(\partial{\mathbb{M}})}{\sqrt{2}\epsilon^{d-1}}\leq k(\epsilon)\leq\frac{1}{\epsilon^{d}}$ where $A(.)$ denotes the surface area in $d-$dimensional space. The lower bound is due to the fact that the area of any slice of the unit hypercube being at-most $\sqrt{2}$ (Ball, 1986). Upper bound is reached when the decision boundary traverses through all the cells in the grid which is less likely in practice. When the model complexity increases, we get a larger $k(\epsilon)$ as well as a smaller $v^{*}(\epsilon)$, requiring a higher number of queries to achieve similar probabilities of success.
Corollary 3.8 (Linear Models).
For linear models with one-sided counterfactuals, ${\mathbb{P}}\left[\text{Reconstruction}\right]=1-(1-v)^{n}$ where $v$ is the volume of the unfavorable region. However, with two-sided counterfactuals, ${\mathbb{P}}\left[\text{Reconstruction}\right]=1$ with just one single query.
This result mathematically demonstrates that allowing counterfactuals from both accepted and rejected regions (as in Aïvodji et al. (2020); Wang et al. (2022)) is easier for model reconstruction, when compared to the one-sided case. It effectively increases each $v_{i}(\epsilon)$ (volume of the inverse counterfactual region). As everything else remains unaffected, for a given $n$, ${\mathbb{P}}[\text{Reconstruction}]$ is higher when counterfactuals from both regions are available. For a linear model, this translates to a guaranteed reconstruction with a single query since $v=1$.
However, we note that all of the aforementioned analysis relies on the closest counterfactual which can be challenging to generate. Practical counterfactual generating mechanisms usually provide counterfactuals that are reasonably close but may not be exactly the closest. This motivates us to now propose a more general strategy assuming local Lipschitz continuity of the models involved.
3.3 Beyond closest counterfactuals
Lipschitz continuity, a property that is often encountered in related works (Bartlett et al., 2017; Gouk et al., 2021; Pauli et al., 2021; Hamman et al., 2023, 2024; Liu et al., 2020; Marques-Silva et al., 2021), demands the model output does not change too fast. Usually, a smaller Lipschitz constant is indicative of a higher generalizability of a model (Gouk et al., 2021; Pauli et al., 2021).
Definition 3.9 (Local Lipschitz Continuity).
A model $m$ is said to be locally Lipschitz continuous if for every $\boldsymbol{x}_{1}\in[0,1]^{d}$ there exists a neighborhood ${\mathbb{B}}_{\boldsymbol{x}_{1}}\subseteq[0,1]^{d}$ around $\boldsymbol{x}_{1}$ such that for all $\boldsymbol{x}_{2}\in{\mathbb{B}}_{\boldsymbol{x}_{1}}$, $|m(\boldsymbol{x}_{1})-m(\boldsymbol{x}_{2})|\leq\gamma||\boldsymbol{x}_{1}-\boldsymbol{x}_{2}||_{2}$ for some $\gamma\in{\mathbb{R}}_{0}^{+}$.
For analyzing model reconstruction under local-Lipschitz assumptions, we consider the difference of the model output probabilities (before thresholding) as a measure of similarity between the target and surrogate models because it would force the decision boundaries to overlap. The proposed strategy is motivated from the following observation: the difference of two models’ output probabilities corresponding to a given input instance $\boldsymbol{x}$ can be bounded by having another point with matching outputs in the affinity of the instance $\boldsymbol{x}$. This observation is formally stated in Theorem 3.10. See Appendix A.5 for a proof.
Theorem 3.10.
Let the target $m$ and surrogate $\tilde{m}$ be ReLU classifiers such that $m(\boldsymbol{w})=\tilde{m}(\boldsymbol{w})$ for every counterfactual $\boldsymbol{w}$. For any point $\boldsymbol{x}$ that lies in a decision boundary cell, $|\tilde{m}(\boldsymbol{x})-m(\boldsymbol{x})|\leq\sqrt{d}(\gamma_{m}+\gamma_{\tilde{m}})\epsilon$ holds with probability $p\geq 1-k(\epsilon)(1-v^{*}(\epsilon))^{n}$.
Note that within each decision boundary cell, models are affine and hence locally Lipschitz for some $\gamma_{m},\gamma_{\tilde{m}}\in{\mathbb{R}}^{+}_{0}$. Local Lipschitz property assures that the approximation is quite close ($\gamma_{m},\gamma_{\tilde{m}}$ are small) except over a few small ill-behaved regions of the decision boundary. This result can be extended to any locally Lipschitz pair of models as stated in following corollary.
Corollary 3.11.
Suppose the target $m$ and surrogate $\tilde{m}$ are locally Lipschitz (not necessarily ReLU) such that $m(\boldsymbol{w})=\tilde{m}(\boldsymbol{w})$ for every counterfactual $\boldsymbol{w}$. Assume the counterfactuals are well-spaced out and forms a $\delta$-cover over the decision boundary. Then $|\tilde{m}(\boldsymbol{x})-m(\boldsymbol{x})|\leq(\gamma_{m}+\gamma_{\tilde{m}})\delta,$ over the target decision boundary.
Figure 6: Rationale for Counterfactual Clamping Strategy.
@@ 3.3 Beyond closest counterfactuals
Theorem 3.10 provides the motivation for a novel model reconstruction strategy. Let $\boldsymbol{w}$ be a counterfactual. Recall that $\partial{\mathbb{M}}$ denotes the decision boundary of $m$. As implied by the theorem, for any $\boldsymbol{x}\in\partial{\mathbb{M}}$, the deviation of the surrogate model output from the target model output can be bounded above by $\sqrt{d}(\gamma_{m}+\gamma_{\tilde{m}})\epsilon$ given that all the counterfactuals satisfy $m(\boldsymbol{w})=\tilde{m}(\boldsymbol{w})$. Knowing that $m(\boldsymbol{w})=0.5$, we may design a loss function which clamps $\tilde{m}(\boldsymbol{w})$ to be 0.5. Consequently, with a sufficient number of well-spaced counterfactuals to cover $\partial{\mathbb{M}}$, we may achieve arbitrarily small $|\tilde{m}(\boldsymbol{x})-m(\boldsymbol{x})|$ at the decision boundary of $m$ (Fig. 6).
@@
However, we note that simply forcing $\tilde{m}(\boldsymbol{w})$ to be exactly equal to 0.5 is quite unstable since in practice the target model’s output $m(\boldsymbol{w})$ is known to be close to 0.5, while being greater but may not be exactly equal. Thus, we propose a unique loss function for training the surrogate neural networks that does not penalize the counterfactuals that are already inside the favorable region of the surrogate model. For $0<k\leq 1$, we define the Counterfactual Clamping loss function as
$\!L_{k}$ $(\tilde{m}(\boldsymbol{x}),y_{\boldsymbol{x}})=\mathds{1}\left[y_{\boldsymbol{x}}=0.5,\tilde{m}(\boldsymbol{x})\leq k\right]\left\{L(\tilde{m}(\boldsymbol{x}),k)-h(k)\right\}+\mathds{1}\left[y_{\boldsymbol{x}}\neq 0.5\right]L(\tilde{m}(\boldsymbol{x}),y_{\boldsymbol{x}}).$ (3)
@@
Here, $y_{\boldsymbol{x}}$ denotes the label assigned to the input instance $\boldsymbol{x}$ by the target model, received from the API. $L(\hat{y},y)$ is the binary cross-entropy loss and $h(\cdot)$ denotes the binary entropy function. We assume that the counterfactuals are distinguishable from the ordinary instances, and assign them a label of $y_{\boldsymbol{x}}=0.5$. The first term accounts for the counterfactuals, where they are assigned a non-zero loss if the surrogate model’s prediction is below $k$. This term ensures $\tilde{m}(\boldsymbol{w})\gtrsim k$ for the counterfactuals $\boldsymbol{w}$ while not penalizing the ones that lie farther inside the favorable region. The second term is the ordinary binary cross-entropy loss, which becomes non-zero only for ordinary query instances. Note that substituting $k=1$ in $L_{k}(\tilde{m}(\boldsymbol{x}),y_{\boldsymbol{x}})$ yields the ordinary binary cross entropy loss. Algorithm 1 summarizes the proposed strategy.
Algorithm 1 Counterfactual Clamping
@@
3:     Query API with $\boldsymbol{x}$ to get $y_{\boldsymbol{x}}$$y_{\boldsymbol{x}}\in\{0,1\}$}
4:     ${\mathbb{A}}\leftarrow{\mathbb{A}}\cup\{(\boldsymbol{x},y_{\boldsymbol{x}}\}$
5:     if $y_{\boldsymbol{x}}=0$ then
@@
10:  Train $\tilde{m}$ on ${\mathbb{A}}$ with $L_{k}(\tilde{m}(\boldsymbol{x}),y_{\boldsymbol{x}})$ as the loss
@@
It is noteworthy that this approach is different from the broad area of soft-label learning Nguyen et al. (2011a, b) in two major aspects: (i) the labels in our problem do not smoothly span the interval [0,1] – instead they are either 0, 1 or 0.5; (ii) labels of counterfactuals do not indicate a class probability; even though a label of 0.5 is assigned, there can be counterfactuals that are well within the surrogate decision boundary which should not be penalized. Nonetheless, we also perform ablation studies where we compare the performance of our proposed loss function with another potential loss which simply forces $\tilde{m}(\boldsymbol{w})$ to be exactly $0.5$, inspired from soft-label learning (see Appendix B.2.10 for results).
Counterfactual Clamping overcomes two challenges beset in existing works; (i) the problem of decision boundary shift (particularly with one-sided counterfactuals) present in the method suggested by Aïvodji et al. (2020) and (ii) the need for counterfactuals from both sides of the decision boundary in the methods of Aïvodji et al. (2020) and Wang et al. (2022).
4 Experiments
We carry out a number of experiments to study the performance of our proposed strategy Counterfactual Clamping. We include some results here and provide further details in Appendix B.
All the classifiers are neural networks unless specified otherwise and their decision boundaries are not necessarily convex. The performance of our strategy is compared with the existing attack presented in Aïvodji et al. (2020) that we refer to as “Baseline”, for the case of one-sided counterfactuals. As the initial counterfactual generating method, we use an implementation of the Minimum Cost Counterfactuals (MCCF) by Wachter et al. (2017).
A summary of the experiments is provided below with additional details in Appendix B.
(i) Visualizing the attack using synthetic data: First, the effect of the proposed loss function in mitigating the decision boundary shift is observed over a 2-D synthetic dataset. Fig. 7 presents the results. In the figure, it is clearly visible that the Baseline model is affected by a decision boundary shift. In contrast, the CCA model’s decision boundary closely approximates the target decision boundary. See Appendix B.2.1 for more details.
Figure 7: A 2-D demonstration of the proposed strategy. Orange and blue shades denote the favorable and unfavorable decision regions of each model. Circles denote the target model’s training data.
(ii) Comparing performance over four real-world dataset: We use four publicly available real-world datasets namely, Adult Income, COMPAS, DCCC, and HELOC (see Appendix B.1) for our experiments. Table 1 provides some of the results over four real-world datasets. We refer to Appendix B.2.2 (specifically Fig. 11) for additional results. In all cases, we observe that CCA has either better or similar fidelity as compared to Baseline.
Table 1: Average fidelity achieved with 400 queries on the real-world datasets over an ensemble of size 100. Target model has hidden layers with neurons (20,10). Model 0 is similar to the target model in architecture. Model 1 has hidden layers with neurons (20, 10, 5).